### v1

# PSR J1922+37: a 1.9-second pulsar discovered in the direction of the old open cluster NGC 6791

Abstract
More than 300 pulsars have been discovered in Galactic globular clusters; however, none have been found in open clusters. Here we present results from 20-hour pulsar searching observations in seven open clusters with the Five-hundred-meter Aperture Spherical radio Telescope (FAST). Our first discovery is a 1.9-second pulsar (J1922+37) found in the direction of the old open cluster NGC 6791. The measured dispersion measure (DM) implies a distance of 4.79 kpc and 8.92 kpc based on the NE2001 and YMW16 electron density models, respectively. Given the large uncertainty of DM distance estimates, it is likely that PSR J1922+37 is indeed a member of NGC 6791, for which the distance is $4.19\pm 0.02$ kpc based on Gaia Data Release 3. If confirmed, PSR J1922+37 will be the first pulsar found in Galactic open clusters. We outline future observations that can confirm this pulsar-open cluster association and discuss the general prospects of finding pulsars in open clusters.
Radio pulsars (1353) — Open star clusters (1160) — Globular star clusters (656)
facilities: FAST, OzStarsoftware: astropy (Astropy Collaboration et al., 2013, 2018), pdmp (Hotan et al., 2004), peasoup (Barr, 2020), psrchive (Hotan et al., 2004)
1 Introduction
Pulsars are well known for their fast spin, high density, and strong gravitational and magnetic field, rendering them astrophysical laboratories unachievable on Earth (see e.g. Lorimer & Kramer, 2012). Globular clusters have the highest spatial concentrations of pulsars, with more than 300 pulsars discovered in 45 globular clusters [^1] https://www3.mpifr-bonn.mpg.de/staff/pfreire/GCpsr.html. In particular, 47 Tucanae alone has more than 40 pulsars, while Terzan 5 has nearly 50 pulsars. The large number of pulsars has provided a unique tool to probe the density profile and the dynamics of their host clusters, and to constrain the mass of potential intermediate massive black holes inside (Freire et al., 2017; Prager et al., 2017; Perera et al., 2017).
The high yield of pulsars in globular clusters is believed to be caused by their high stellar encounter rate, which is directly related to their crowded stellar environments (Hui et al., 2010; Bahramian et al., 2013). Specifically, the encounter rate, $\Gamma$, is proportional to $\rho_{\rm c}^{2}r_{\rm c}^{3}/v$, where $\rho_{\rm c}$ and $r_{\rm c}$ are the density and size of the core, respectively, and $v$ is the velocity dispersion of the cluster (Verbunt, 2002; Verbunt & Freire, 2014), while the number of pulsar discoveries, $\lambda$, is expected to follow the scaling relation $\log\lambda=-1.1+1.5\log_{10}\Gamma$ (Turk & Lorimer, 2013). More recently, Yin et al. (2024) showed that escape velocity could also be a determining factor for $\lambda$.
Open clusters also have a relatively high stellar density, and more than 2900 such clusters have been identified, about 10 times that of globular clusters (e.g. Bica et al., 2019, Table 1). However, no pulsar discoveries have been confirmed in any open clusters yet. The difference is probably caused by the lower stellar density and thus a lower encounter rate in open clusters. In addition, the lower escape velocity of open clusters also makes it more difficult to retain pulsars that usually suffer from natal kicks (e.g. Pfahl et al., 2002; Ivanova et al., 2008; Boyles et al., 2011). The short lifetime ($\sim 10^{8}$ yr) of open clusters also means that they are likely to have disintegrated before any pulsars were discovered inside. As a result, open clusters were unfavorable places for pulsar searching and were usually overlooked in past search efforts. However, in the search for unusual types of pulsar, open clusters deserve more attention. Any confirmed pulsar discovery in these clusters will greatly advance our understanding of stellar evolution and pulsar formation.
As the largest single-dish telescope, the Five-hundred-meter Aperture Spherical radio Telescope (FAST) (Nan et al., 2011) provides an unprecedented sensitivity for pulsar searches. In addition, the adoption of a 19-beam receiver also provides a decent speed to observe a large sky area (Jiang et al., 2020). Thus, FAST has been proven to be a productive telescope for pulsar discovery. To date, FAST has discovered more than 800 new pulsars in the past few years, mainly through the Galatic Plane Pulsar Survey (GPPS, Han et al. 2021), the Commensal Radio Astronomy FAST Survey (CRAFTS, Li et al. 2018) and dedicated globular cluster searches (e.g. Pan et al., 2021).
In this paper, we present results from a 20-h FAST observing project (ID: PT2021_0102), with the goal of finding exotic pulsar systems in seven open clusters. After initial analysis of these observations, we found a 1.9-second pulsar, J1922+37, which is likely to be associated with open cluster NGC 6791. If confirmed, J1922+37 will be the first pulsar discovered in an open cluster and will provide new insights into the stellar evolution in open clusters.
The remainder of the paper is organized as follows. We first describe the observations in Section 2, then introduce the search method in Section 3 and present our results in Section 4, before discussing the results in Section 5. The paper is concluded in Section 6.
2 Observations
Using the Milky Way star cluster catalog [^2] https://heasarc.gsfc.nasa.gov/W3Browse/star-catalog/mwsc.html (Kharchenko et al., 2013), we chose seven open clusters that are old and massive as targets. The basic information of the clusters is shown in Table 1. We see that all clusters are within $\sim 10^{\circ}$ of the Galactic plane. If no pulsars are found in the open clusters, we may still be able to detect pulsars in the foreground or background sky; note that the same sky locations (of these open clusters) have been or will be covered by the GPPS survey (Han et al., 2021), which motivates us to adopt a longer integration time for our observations (see Tables 2) and 3 for details).
Observations were carried out using all beams of the L-band 19-beam receiver (Jiang et al., 2020) with a time resolution of 49.152 $\mu$s. Two polarizations and 4096 frequency channels were used throughout the observations. The receiver provides a bandwidth of 500 MHz, but only 400 MHz are effective. Data were recorded with the pulsar backend in search mode and saved in psrfits format (Hotan et al., 2004) with a default depth of 8 bits. For calibration purposes, each observation began with a low-intensity noise injection, which has a modulating period of 0.2 second and lasted for one minute.
Taking into account the gaps between individual beams, four pointings are necessary to cover a given region (e.g. Jiang et al., 2020). For efficient coverage, all clusters were observed using two groups of on-off mode observations [^3] https://fast.bao.ac.cn/cms/article/24/, except for IC 4756, which was observed with the more efficient snapshot mode, the same observation mode used for GPPS (Han et al., 2021). Due to the large angular size of IC 4756, two groups of snapshot-mode observations were applied to cover the central region of the cluster. The observation position and integration length of the clusters are given in Table 2 (for the on-off mode) and Table 3 (for the snapshot mode). For the on-off mode, we provide the central beam (M01) of each pointing, while for the snapshot mode, only the central beam of the first pointing is provided. In both modes, the position of the remaining beams can be deduced by accounting for the geometry of the 19-beam receiver and considering the slewing track of the receiver in each mode (Jiang et al., 2020; Han et al., 2021).
Table 1: Summary of the seven open clusters observed in this project. Columns are the cluster name, equatorial and Galactic coordinates, tidal radius, age, distance, and the estimated dispersion measure (DM) based on the distance and the NE2001 and YMW16 electron density model. The cluster parameters have been updated using Hunt & Reffert (2024).
Open cluster RA DEC Gl Gb $R$ Age $D$ DMNE2001 DMYMW16
(hhmmss) (ddmmss) (deg) (deg) (arcmin) ($10^{8}$ yr) (kpc) (cm-3 pc) (cm-3 pc)
NGC 1912/M38 05:28:40 +35:49:15 172.27 0.68 40.4 2.7 1.09 34.8 38.5
NGC 2099/M37 05:52:19 +32:32:27 177.65 3.09 30.4 4.9 1.40 46.5 60.3
NGC 2158 06:07:26 +24:05:52 186.64 1.78 12.6 18.6 3.71 128.2 172.8
NGC 2168/M35 06:08:48 +24:17:11 186.62 2.15 14.3 6.3 2.40 91.2 149.5
IC 4756 18:38:31 +05:26:14 36.31 5.34 94.6 8.4 0.47 4.8 7.3
NGC 6791 19:20:53 +37:46:27 69.96 10.91 11.6 38.0 4.19 74.1 49.6
NGC 7789 23:57:19 +56:43:26 115.52 $-5.37$ 32.2 15.7 1.97 44.2 47.9
Table 2: The observation position, integration length and observation date of open clusters observed with the on-off mode. Each cluster was observed with two groups of on-off mode hence four pointings. For each pointing, the position of the central beam (M01) is listed. $t_{\rm int}$ is the approximate integration length of each pointing, while the observation date is based on UTC+8.
Open cluster on RA on DEC off RA off DEC $t_{\rm int}$ Obs. Date
(hhmmss) (ddmmss) (hhmmss) (ddmmss) (min) (yy-mm-dd)
NGC 1912/M38 05:28:52.00 +35:48:00.0 05:29:06.80 +35:48:00.0 27 2021-09-27
05:28:44.60 +35:50:36.0 05:28:59.40 +35:50:36.0
NGC 2099/M37 05:52:21.00 +32:34:12.0 05:52:35.24 +32:34:12.0 27 2021-09-28
05:52:13.88 +32:36:48.0 05:52:28.12 +32:36:48.0
NGC 2158 06:07:26.00 +24:05:30.0 06:07:39.15 +24:05:30.0 57 2021-09-29
06:07:19.43 +24:08:06.0 06:07:32.57 +24:08:06.0
NGC 2168/M35 06:09:13.00 +24:21:36.0 06:09:26.17 +24:21:36.0 27 2021-09-30
06:09:06.41 +24:24:12.0 06:09:19.59 +24:24:12.0
NGC 6791 19:20:53.00 +37:46:48.0 19:21:08.18 +37:46:48.0 57 2022-01-22
19:20:45.41 +37:49:24.0 19:21:00.59 +37:49:24.0
NGC 7789 23:57:25.00 +56:43:48.0 23:57:46.87 +56:43:48.0 27 2021-09-29
23:57:14.06 +56:46:24.0 23:57:35.94 +56:46:24.0
Table 3: The observation position, integration length and observation date of IC 4756. Since the snapshot mode was used for the cluster, only the position of the central beam (M01) in the first pointing is shown.
Open cluster RA DEC $t_{\rm int}$ Obs. Date
(hhmmss) (ddmm) (min) (yy-mm-dd)
IC 4756 18:39:38.19 +05:29:36.0 18 2022-01-29
18:38:01.75 +05:29:36.0 28 2022-02-03
3 Search method
A bottleneck of processing pulsar search mode data is the slow I/O speed due to the large data volume. To reduce data size, we convert the data from 8 bits to 2 bits and halve the number of frequency channels using the digifil routine in the dspsr software package (van Straten & Bailes, 2011). The conversion thus reduces the size of the data by a factor of 8, but only leads to a few percent decrease in signal-noise-ratio (S/N), as demonstrated using the known pulsar J1837+0528g [^4] http://zmtt.bao.ac.cn/GPPS/GPPSnewPSR.html in the line-of-sight of IC 4756 (see Section 4.1). After conversion, the data have 2048 frequency channels while the time resolution remains unchanged.
To minimize the impact of radio frequency interference (RFI), we found the bad frequency channels in each observation using the rfifind routine in prestosoftware package (Ransom, 2011). A length of one second was used to integrate for the analysis. The bad frequency channels were masked out for the remaining process.
We also removed the periodic interference using the sigpyproc software package [^5] https://sigpyproc3.readthedocs.io/en/latest/. For this purpose, we first removed the bad frequency channels from each observation. For each beam, a time series was then generated at 0 DM and a Fast Fourier Transform (FFT) was performed to obtain the power spectrum using sigpyproc. Since the receiver’s 19 beams observe the sky simultaneously, any periodic interference would appear across multiple beams. Therefore, to identify such periodic radio frequency interference or “birdies”, we flagged frequency bins if they exceeded a certain threshold ($\sigma>4$) in the power spectrum of four or more beams. These frequencies were used as a common birdie list for all 19-beams of a single pointing.
To de-disperse the data, we searched the range of DM in [2, 600] cm-3 pc, where the upper end is much larger than the largest DM value estimated using the cluster distance and the NE2001 (Cordes & Lazio, 2002) or YMW16 (Yao et al., 2017) electron density model, see Table 1.
To minimize the loss of sensitivity to pulsars in compact binaries (e.g. Johnston & Kulkarni, 1991; Ransom et al., 2003), we also searched in acceleration space. Considering the computational cost and the large number of candidates for a thorough search, we first conducted a fast search in a small acceleration range, $|a|\leq 1$ m s-2, while a deeper search considering a larger acceleration range is ongoing and the results will be presented elsewhere.
Both dedispersion and acceleration search were performed with the peasoup [^6] https://github.com/ewanbarr/peasoup software package, which can also take into account the bad frequency channels and birdies mentioned above. More importantly, peasoup is based on the Graphic Processing Unit (GPU), which can significantly speed up the search process (e.g. Morello et al., 2019; Sengar et al., 2023). For each search, in spectral domain, we set an S/N threshold of 6.5 and only kept the candidates above the threshold. As a result, 15000 candidates were found for each pointing.
To reduce the number of candidates to a manageable scale, we adopted a parameter-based method, which uses a combination of the candidate period, S/N and the number of harmonics to pick out the most probable pulsar candidates (Sengar et al., 2023). We then folded the search mode data using the period, DM, and acceleration, with the dspsr routine in the psrchive [^7] https://psrchive.sourceforge.net/ software package (Hotan et al., 2004). The resulting archive files were first cleaned with the interference removal program clfd [^8] https://github.com/itachi-gf/clfd/tree/master (Morello et al., 2023), before being processed with the pdmp routine in psrchive, which optimizes the period, DM, and S/N and provides diagnostic plots for the inspection. Using the refined parameters, the sifting and folding process was repeated once, further reducing the number of candidates and creating diagnostic plots for these candidates, which were visually inspected to pick out pulsars and highly confident candidates.
4 Results
4.1 Discovery and confirmation
Two pulsars were found in the search, PSRs J1837+05 and J1922+37. PSR J1837+05 is a 6.26 ms pulsar and has a DM of $\sim 120.8$ pc cm-3, consistent with PSR J1837+0528g, which was discovered by GPPS. J1837+05 was detected in three adjacent beams, which are also consistent with the discovery position of J1837+0528g, confirming the re-detection nature. J1837+05 is in the line-of-sight of open cluster IC 4756, but its DM is about 20 times the DM estimate using the cluster distance and the electron density models (see Table 1). J1837+05 is thus more likely in the background than in the open cluster.
PSR J1922+37 has a period of 1.92 s and a DM of $\sim 85$ pc cm-3 (Fig. 1), and is in the line-of-sight of open cluster NGC 6791 (Fig. 2). The pulsar was discovered on 22 January 2022 in two adjacent beams (M08) with a comparable S/N $\sim 60$. In addition to the fundamental spin period, harmonics at the 2nd, 3rd, 6.5th, 17th, 19th, 21st and 25th were also detected. No significant acceleration was detected. Using FAST, the pulsar was later confirmed in an 18-min follow-up observation and in a 6-min grid observation (see Section 4.3 for more details).
Figure 1: The discovery plot of PSR J1922+37 from the pdmp routine in psrchive. The top panel shows the optimization of period and DM, while the middle panels are phase-time and phase-frequency plots and the bottom panel shows the pulse profile.
Figure 2: The position of J1922+37 and the member stars of NGC 6791. The filled circles in orange, pink, cyan, and olive show the tiled beams of all four observation pointings. The two discovery beams (M08) are indicated by the hollow stars, while the optimized position is labeled by the solid red star, and the cluster members are marked by the gray stars. The blue and red circles show the boundary of the tidal radius and Jacobi radius, respectively.
4.2 Pulse width and flux density
To estimate the flux density of PSR J1922+37, we measured its pulse width. We fitted the pulse region with a two-component Gaussian model to account for both the main- and sub-pulse. As an approximation, we use the full width at half maximum (FWHM) of the main pulse as the pulse width, $0.012\pm 0.002$ s, which gives a duty cycle of $0.63\%\pm 0.08\%$.
We estimate the flux density by the radiometer equation (Lorimer & Kramer, 2012),
$S=\frac{{\rm S/N}~{}T_{\rm sys}}{G\sqrt{n_{\rm p}t_{\rm int}\Delta f}}\sqrt{\frac{W}{P-W}},$ (1)
where we use $T_{\rm sys}=18$ K, $\Delta f=400$ MHz (Jiang et al., 2019), and $G$ is the effective gain. For the discovery observation, which has the highest S/N = 66, $n_{\rm p}=2$ and $t_{\rm int}=3400$ s. According to Jiang et al. (2020), the gain of beam M08 is about 0.92 times that of the central beam at 1250 MHz, which gives $G_{0}=14.8$ K Jy-1. Note that the center of the discovery beams is about half a beam-width off the source, so the effective gain should be smaller. Assume a Gaussian beam pattern (e.g. Condon & Ransom, 2016), then $G=G_{0}/2$. Using the parameters above, we obtain a flux density of $\sim 7~{}\mu$Jy.
All the observations mentioned above were spread over 3 years and thus are insufficient to produce a phase-coherent timing solution. As a result, spin-down and characteristic age are unavailable at the moment. No reliable polarization information was obtained, as all the observations either do not have sufficiently high S/N or full Stokes parameters, making the polarization measurement insignificant.
4.3 Position and distance
We refined the pulsar position using two groups of on-off-mode observations with FAST and the same receiver. The observations were centered on the midpoint of the two adjacent discovery beams and fully covered the nearby region in grids. Each pointing was 6 minutes. The pulsar was detected in three adjacent beams, one beam reporting an S/N of $\sim 17$ and two beams of $\sim 7$. Using the S/N-weighted sum of the center of the beams, we update the pulsar position to RA = 19:22:03 and Dec = +37:46:13 with an uncertainty of the beam size ($\sim 3\arcmin$).
According to the position, PSR J1922$+$37 is in the same direction as NGC 6791 with an offset of $\sim 14\arcmin$ from the center of the cluster. Fig. 2 shows the member stars of NGC 6791, which are taken from Hunt & Reffert (2024) and have a membership probability greater than 30% [^9] https://cdsarc.cds.unistra.fr/viz-bin/cat/J/A+A/686/A42#/article. PSR J1922+37 is slightly outside the King (1962) tidal radius of NGC 6791 but still within its Jacobi radius, an alternative to estimate the tidal radius (Binney & Tremaine, 2008; Hunt & Reffert, 2024).
Moreover, using the optimal DM and position, the NE2001 electron density model (Cordes & Lazio, 2002) gives a distance estimate of $d_{\rm NE2001}\sim 4.79$ kpc, which is close to the cluster distance reported by the Milky Way star cluster survey (4.93 kpc, Kharchenko et al. 2013) and that inferred from Gaia Data Release 2 ($4.45\pm 0.10$ kpc, Dias et al., 2021), but slightly differs from that obtained using red clump giants in the cluster ($4.02\pm 0.15$ kpc Gao & Chen, 2012) and that using Gaia Data Release 3 ($4.19\pm 0.02$ kpc, Hunt & Reffert, 2024). However, the distance estimated by the YMW16 electron density model (Yao et al., 2017) is $d_{\rm YMW16}\sim 8.92$ kpc, about two times the distance to NGC 6791.
5 Discussion
The small angular separation and the close distance predicted by the NE2001 model make PSR J1922+37 a likely member of NGC 6791. Actually, three Globular cluster pulsars are significantly far from corresponding cluster center, including PSRs B1718--19A (NGC 6342, Lyne et al., 1993), J1911--5958A (NGC 6752, D’Amico et al., 2001) and J1801--0857D (NGC 6517, Lynch et al., 2011). A fourth pulsar, J1823--3022, is at $\sim 3$ times the half-light radius of NGC 6624, although their association still awaits confirmation (Abbate et al., 2022).
If the association is confirmed, PSR J1922+37 will be the first pulsar found in an open cluster, and more pulsars may be discovered in NGC 6791. However, the large discrepancy between $d_{\rm YMW16}$ and the distance to NGC 6791 suggests that a coincidental overlap is also possible. Considering the large uncertainty of distance estimates from electron density models, it is vital to check the association with further investigations.
There are several ways to check if PSR J1922+37 is actually associated with NGC 6791. First, a precise measurement of the distance to the pulsar and the open cluster will give a straightforward answer to the problem. The recent Gaia data release has allowed a much better estimation of the distance to NGC 6791 (e.g. Dias et al., 2021; Hunt & Reffert, 2024), and the estimate will likely improve in future data releases. However, the low flux density of PSR J1922+37 will pose a challenge to measuring its distance with interferometry. The large distance also makes measurement with timing parallax difficult.
Second, if PSR J1922+37 is associated with NGC 6791, then both objects should have similar proper motion. Bedin et al. (2006) reported a proper motion of $(\mu_{\alpha}\cos\delta,\mu_{\delta})=(-0.57,-2.45)$ mas yr-1 for NGC 6791, while Hunt & Reffert (2024) gave $(-0.42,-2.28)$ mas yr-1. A measurement of the proper motion of PSR J1922+37, e.g. via pulsar timing, will be helpful in determining the relation between the two objects.
Finally, like the pulsars found in globular clusters, if more pulsars with a similar DM can be found in the direction of NGC 6791, then there is a high probability that PSR J1922+37 is associated with the open cluster.
5.1 A pulsar population in NGC 6791?
NGC 6791 is an unusually old and massive open cluster in our Galaxy (Grundahl et al., 2008), with an age of several Gyr (Dias et al., 2021; Hunt & Reffert, 2024) and a total mass ranging from 5000 M (Platais et al., 2011) to a few $10^{4}$ M (Hunt & Reffert, 2024). The cluster is relatively compact and follows a single-mass King density profile up to $\sim 600\arcsec$, with a core region of $r_{\rm c}\sim 2\farcm 66$, compactness of $c\sim 1.06$ and a central surface brightness of $\mu\sim 17.5$ mag per asrcsec2 (Dalessandro et al., 2015). The relatively high surface brightness and compactness thus lead to a crowded stellar environment, making NGC 6791 a possible birthplace of pulsars, similar to the pulsars in globular clusters (Verbunt, 2002; Hui et al., 2010; Bahramian et al., 2013).
Following Bahramian et al. (2013), we estimate the expected number of pulsar discoveries, $\lambda$, in NGC 6791 in three steps. First, using the cluster parameters above, we computed the central luminosity density, $\rho_{\rm c}$, according to the prescription given by Djorgovski (1993). Second, we obtained the encounter rate $\Gamma\sim\rho_{\rm c}^{2}r_{\rm c}^{3}/\sigma_{v}\approx 162$ assuming a King potential model (Dalessandro et al., 2015) and a velocity dispersion of $\sigma_{v}\sim 2$ km s-1. Note that $\Gamma$ has been normalized so that the encounter rate of 47 Tuc is $\Gamma_{\rm 47~{}Tuc}=1000$ (Bahramian et al., 2013). Finally, the expected number of pulsars is found by $\ln\lambda=-1.1+1.5\log_{10}\Gamma$ (Turk & Lorimer, 2013; Gao et al., 2024), which gives $\lambda\sim 9$ potential discoveries. This result is also in general agreement with the estimation using escape velocity; see Yin et al. (2024, Fig. 2). Therefore, more pulsars are likely in NGC 6791, and it deserves future observations.
5.2 Potential pulsar discoveries in other open clusters
The possible association motivates us to check if there are other open clusters that may also host pulsars. A detailed analysis inspecting the association between all known open clusters and pulsars is beyond the scope of this work, and will be presented in Zhou et al. (2024, in preparation). Here, we check the latest Galactic cluster catalog (Hunt & Reffert, 2024) and highlight the open clusters that are worth a careful pulsar search.
We cannot obtain $\lambda$ for each open cluster due to the unavailability of $\rho_{\rm c}$. Instead, we sift through the catalog and pick out the open clusters that are similar to NGC 6791 in both total mass ($>10^{4}$ M) and core size ($<10$ pc), selecting open clusters with comparable core density and thus encounter rate. We also set a lower age limit of 100 Myr, which is a typical lifetime for canonical pulsars.
Ten open clusters, including NGC 6791, meet the above conditions and are shown in Table 4, which lists the total mass, radial velocity scattering, angular size, and number of member stars. All the clusters are relatively close ($d<5$ kpc) and have a small core size with a typical value of $\sim 5\arcmin$. They also have a relatively high star concentration, with half of members within the tidal radius. These clusters are thus good candidates for future pulsar search in open clusters.
[…]
(deg) (deg) (kpc) ($10^{8}$ yr) ($10^{4}$ M) (km s-1) (pc) (min) (min) (min)
Berkeley 53 313.98 51.05 3.41 5.6 3.73 0.6 3.9 3.9 11.4 26.7 714 1660
Theia 1661 267.59 $-25.15$ 2.30 2.4 2.71 2.6 2.6 3.9 105.5 113.9 2607 2620
Trumpler 5 99.13 9.46 2.90 21.7 2.46 0.6 3.9 4.6 30.2 44.2 2551 3619
NGC 2158 91.86 24.10 3.71 18.6 2.09 1.3 2.0 1.8 12.6 39.6 1207 1958
IC 166 28.09 61.86 4.42 7.8 2.07 0.7 2.7 2.1 11.5 28.5 801 1768
NGC 6791 290.22 37.77 4.19 38.0 1.95 2.2 4.5 3.7 11.6 41.5 2062 4193
NGC 7789 359.33 56.72 1.97 15.7 1.78 0.9 4.0 6.9 32.2 49.5 2377 4125
NGC 6259 255.19 $-44.68$ 2.16 1.6 1.78 1.8 4.1 6.6 23.4 34.7 1310 2188
Pismis 3 127.83 $-38.66$ 2.11 10.6 1.59 0.8 3.6 5.8 22.0 68.1 1204 2879
Trumpler 20 189.89 $-60.63$ 3.28 15.8 1.53 0.5 9.7 10.2 10.2 53.4 811 1692
Table 4: Open clusters similar to NGC 6791 in decreasing total mass. Columns are cluster name, right ascention and declination, distance, age, total mass, the standard error of radial velocity, core radius (in pc and in arcmin), tidal radius, total radius, number of member stars within $r_{\rm tidal}$ and the total number of stars.
6 Conclusions
Pulsars located or born in open clusters have not drawn much attention in pulsar research, but may hold important clues in understanding the formation and evolution of different species of pulsars and star clusters. In this paper, we present initial results from 20-hour pulsar searching observations in seven open clusters using FAST. We discovered a new pulsar J1922+37 in the direction of open cluster NGC 6791 (Fig. 2); it has a period of 1.92 second and a DM of $\sim 85$ pc cm-3. The NE2001 model gives a distance estimate of 4.79 kpc to the pulsar, consistent with the cluster distance, which ranges from 4.02 to 4.93 kpc in the literature. The pulsar is thus a likely member of NGC 6791. If confirmed, PSR J1922+37 will be the first pulsar discovered in an open cluster.
Measurements of pulsar distance and proper motion will be crucial in determining whether PSR J1922+37 is associated with NGC 6791 or not. Assuming that old and massive open clusters such as NGC 6791 are similar to Galactic globular clusters as pulsar birth places (e.g., the size of pulsar population is proportional to the stellar encounter rate), we estimate that around 9 pulsars are expected to be discovered in NGC 6791. We also highlight 10 open clusters that are promising sites for finding new pulsars.
Follow-up observations have been planned with FAST, which will allow measurements of spin and astrometric parameters of PSR J1922+37 as well as a deeper search for more pulsars in NGC 6791. More pulsar discoveries with similar DM to PSR J1922+37 in the region will provide compelling evidence for pulsars in open clusters.
We thank Lei Zhang, Pei Wang, Zhichen Pan, Lei Qian and Zucheng Chen for helpful discussions. This work made use of the data from FAST (Five-hundred-meter Aperture Spherical radio Telescope) (https://cstr.cn/31116.02.FAST). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. The data processing was mainly performed on the OzSTAR national facility at Swinburne University of Technology. The OzSTAR program receives funding in part from the Astronomy National Collaborative Research Infrastructure Strategy (NCRIS) allocation provided by the Australian Government, and from the Victorian Higher Education State Investment Fund (VHESIF) provided by the Victorian Government. R.P.E. is supported by the Chinese Academy of Sciences President’s International Fellowship Initiative, grant No. 2021FSM0004. JPY, NW, WWZ are supported by the National Natural Science Foundation of China (grant No. 12041304, 12041303). WWZ is supported by National SKA Program of China (grant No. 2020SKA0120200). XJZ is supported by the National Natural Science Foundation of China (Grant No. 12203004) and by the Fundamental Research Funds for the Central Universities.